# Reconstructing Physics-Informed Machine Learning for Traffic Flow Modeling: a Multi-Gradient Descent and Pareto Learning Approach

### v1

@@ Abstract
Physics-informed machine learning (PIML) is crucial in modern traffic flow modeling because it combines the benefits of both physics-based and data-driven approaches. In conventional PIML, physical information is typically incorporated by constructing a hybrid loss function that combines data-driven loss and physics loss through linear scalarization. The goal is to find a trade-off between these two objectives to improve the accuracy of model predictions. However, from a mathematical perspective, linear scalarization is limited to identifying only the convex region of the Pareto front, as it treats data-driven and physics losses as separate objectives. Given that most PIML loss functions are non-convex, linear scalarization restricts the achievable trade-off solutions. Moreover, tuning the weighting coefficients for the two loss components can be both time-consuming and computationally challenging. To address these limitations, this paper introduces a paradigm shift in PIML by reformulating the training process as a multi-objective optimization problem, treating data-driven loss and physics loss independently. We apply several multi-gradient descent algorithms (MGDAs), including traditional multi-gradient descent (TMGD) and dual cone gradient descent (DCGD), to explore the Pareto front in this multi-objective setting. These methods are evaluated on both macroscopic and microscopic traffic flow models. In the macroscopic case, MGDAs achieved comparable performance to traditional linear scalarization methods. Notably, in the microscopic case, MGDAs significantly outperformed their scalarization-based counterparts, demonstrating the advantages of a multi-objective optimization approach in complex PIML scenarios.
@@ 1 Introduction
PIML effectively combines the advantages of data-driven and physics-based methods, making it useful in transportation research. Data-driven approaches have gained popularity due to their low computational cost and flexibility in managing complex scenarios without requiring strong theoretical assumptions. Many of machine learning techniques, especially deep learning architectures Ma et al. [2015], Duan et al. [2016], Polson and Sokolov [2017], Zhou et al. [2017], Wu et al. [2018], Cui et al. [2018, 2019], have various application in transportation research. However, these models heavily depend on the quality and representativeness of the training data. As shown in Yuan et al. [2021b], when manually adding errors to the training dataset, the performance of all pure data-driven models has significantly deteriorated. While a data-driven approach can yield positive results with quality data, it often lacks interpretability, which may diminish its practicality in decision-making processes.
@@
In nearly all PIML studies, especially in PIML application in transportation, as shown in Table 1, researchers commonly use a method called linear scalarization to define the loss function [^1] In many Physics-regularized Gaussian process models (Yuan et al. [2020],Yuan et al. [2021b],Yuan et al. [2021a]), the training target is to minimize a evidence lower bound, which can be written as: $\log\mathcal{N}([\cdot])+\gamma\mathbb{E}(\cdot)$ (1) where $\boldsymbol{\theta}^{data}$ and $\boldsymbol{\theta}^{physics}$ are exclusive parameters of the data-driven likelihood and the physics regularization, respectively, $\gamma$ is a regularization coefficient. The overall ELBO can also be regarded as linear scalarization., as demonstrated in Eq 2. where $\alpha$ and $\beta$ are hyperparameters that control the effect of the data-driven module and physics module, $\mathcal{L}_{data}$ and $\mathcal{L}_{physics}$ represent data-driven loss term and physics loss term, respectively. This approach facilitates the easy integration of physics information into PIML models. When $\alpha\rightarrow 0$, the PIML model degenerates to a pure data-driven model; when $\beta\rightarrow 0$, the PIML model degenerates to a pure physics-driven model.
@@
In this paper, we aim to circumvent the hyperparameter search process by transforming the training process of the PIML model into a multi-objective optimization problem. Mathematically, we transfer the overall training process of PIML models from Eq 3 to 4.
@@ 2 Literature review
One clear strength of the PIML framework is its flexibility: the learning module can adopt almost any neural-network architecture—or even non-NN techniques—to suit a particular task. For instance, Xue et al. [2024] proposes a Graph Neural Network (GNN) with the Network Macroscopic Fundamental Diagram (NMFD) to impute missing traffic states. The GNN captures complex spatial dependencies among links, while the NMFD term keeps predictions physically plausible. Likewise, Pereira et al. [2022] show that a long short-term memory (LSTM) backbone plugs in just as naturally, enabling short-range traffic forecasting when strong temporal correlations exist. In both cases, the physics term reduces data needs and stabilizes training.
Not all PIML methods rely on neural networks alone. Inspired by Wang et al. [2020], Yuan et al. [2021b] introduce the Physics-Regulated Gaussian Process (PRGP). Here, the governing equations appear through a shadow Gaussian process rather than an extra PINN loss term. Training maximizes a compound evidence lower bound (ELBO) that balances the usual GP likelihood with a penalty that measures how well samples obey the flow equation. A follow-up study by Yuan et al. [2021a] extends this idea to a richer family of macroscopic models, improving fidelity and robustness without sacrificing the analytic uncertainty estimates that make Gaussian processes attractive.
@@ 3.1 Multi-gradient descent algorithms
    There exist $\alpha^{1},...,\alpha^{j}$ such that $\sum_{t=1}^{N}\alpha^{j}=1$ and $\sum_{j=1}^{N}\alpha^{j}\nabla_{\boldsymbol{\theta}^{sh}}\mathcal{L}^{j}(\boldsymbol{\theta}^{sh},\boldsymbol{\theta}^{j})=0$
@@
    For all objectives $j$, $\nabla_{\boldsymbol{\theta}^{j}}\mathcal{L}^{j}(\boldsymbol{\theta}^{sh},\boldsymbol{\theta}^{j})=0$
where $\theta^{sh}$ is the shared parameters among all objectives, $\theta^{j}$ is the exclusive parameters of $j$th objective. [^2] In the context of PIML, $\boldsymbol{\theta}^{sh}\subset\boldsymbol{\theta}^{data}\cap\boldsymbol{\theta}^{physics}$
In addition, readers can find out that some of the studies about multi-gradient descent (Hwang and Lim [2024]) do not divide the total parameter set $\boldsymbol{\theta}$ into $\boldsymbol{\theta}^{sh}$, $\boldsymbol{\theta}^{data}$ and $\boldsymbol{\theta}^{physics}$ like in Sener and Koltun [2018]. We only use $\theta_{t}$ to represent the total parameter set in the $t$th iteration. In fact, let us consider the data-driven loss $\mathcal{L}_{data}$, if we calculate its derivatives of the entire parameter set $\boldsymbol{\theta}$, we will have:
@@ 3.1.1 Traditional multi-gradient descent algorithm
The convex hull of a set of points $S$ in $n$ dimensions is defined as the intersection of all convex sets containing $S$. For $N$ points $\mathbf{p}_{1}$, $\mathbf{p}_{2}$,..,$\mathbf{p}_{N}$, the convex hull $\textbf{conv}(\mathbf{p}_{1},\mathbf{p}_{2},..,\mathbf{p}_{N})$ is then given by:
@@
Let $C\subset\mathbb{R}^{n}$ be a nonempty closed convex set. The minimum‐norm point of $C$ is the unique solution
@@
Let $P$subset $\mathbb{R}^{d}$ be the convex hull of finitely many points, $P$ = $\mathbf{conv}(\mathbf{p}_{1}$, $\mathbf{p}_{2}$,..,$\mathbf{p}_{n}$). A point $\mathbf{x}\in P$ is the minimum norm point in $P$ if and only if:
@@
Let $P$subset $\mathbb{R}^{d}$ be the convex hull of the gradients of the different objective functions, i.e., $P$ = conv($\mathbf{p}_{1}$, $\mathbf{p}_{2}$,..,$\mathbf{p}_{n}$). If $\mathbf{x}$ is the minimum norm point, and $\mathbf{x}\in P$, then $-\mathbf{x}$ is a direction that can improve all objectives.
@@
$-\mathbf{x}^{\top}\mathbf{p}_{j}\leq-\|\mathbf{x}\|_{2}^{2}\leq 0\quad\text{for all }j=1,2,\ldots,n.$ (11)
Recall that $\mathbf{p}_{j}$ is the gradient with respect to $j$th objective function. From (10), the directional derivative of each objective in the direction $-\mathbf{x}$ is non-positive. Therefore, $-\mathbf{x}$ constitutes a direction that does not increase (and thus improve) all objectives. In particular, when $\mathbf{x}$neq $\mathbf{0}$, $-\mathbf{x}$ strictly decreases (and thus improves) all objectives, completing the proof. ∎
@@ 10     $\hat{t}=\operatorname*{arg\,min}_{r}\sum_{t}\alpha^{t}\mathbf{M}_{rt}$
11     $\hat{\gamma}=\operatorname*{arg\,min}_{\gamma}((1-\gamma)\boldsymbol{\alpha}+\gamma\mathbf{e}_{\hat{t}})^{\top}((1-\gamma)\boldsymbol{\alpha}+\gamma\mathbf{e}_{\hat{t}})$
@@
Either solution to this optimization problem is 0 and the resulting point satisfies the KKT conditions, or the solution gives a descent direction that improves all objectives.
@@ 3.1.2 Dual cone gradient descent
[…]
@@
The concept of dual cone shown in Definition 6 plays an important role in the DCGD algorithm. However, in practice, we are more interesting of its subspace $\mathbf{G}_{t}$, as shown in Figure 1 and mathematically defined in Eq 15. Any direction such that $\mathbf{d}\in\mathbf{G}_{t}$ is a common descent direction since $\langle\mathbf{d},\nabla\mathcal{L}_{r}(\theta_{t})\rangle\geq 0$ and $\langle\mathbf{d},\nabla\mathcal{L}_{b}(\theta_{t})\rangle\geq 0$. Also, from Eq 15, $\mathbf{G}_{t}$ is essentially a conic combination of $\nabla\mathcal{L}(\theta_{t})_{\|(\nabla\mathcal{L}_{r}(\theta_{t}))^{\perp}}$ and $\nabla_{t}\mathcal{L}(\theta_{t})_{\|(\nabla\mathcal{L}_{b}(\theta_{t}))^{\perp}}$, which is very easy to compute compare to compute the $\mathbf{K}^{*}_{t}$.
@@
$\mathbf{G}_{t}:=\left\{c_{1}\nabla\mathcal{L}(\theta_{t})_{\|(\nabla\mathcal{L}_{r}(\theta_{t}))^{\perp}}+c_{2}\nabla_{t}\mathcal{L}(\theta_{t})_{\|(\nabla\mathcal{L}_{b}(\theta_{t}))^{\perp}}\big{|}c_{1},c_{2}\geq 0,c_{1},c_{2}\in\mathbb{R}\right\}$ (15)
@@
Algorithm 2 shows the core idea of DCGD, that is to find a common descent direction from $\mathbf{G}^{*}_{t}$, and since any direction that subset to $\mathbf{G}_{t}$ is a valid common descent direction. Thus, as shown in Figure 2, there are several variants of DVGD algorithms [^4] For notation simplicity, $\nabla\mathcal{L}_{\|\nabla\mathcal{L}_{r}^{\perp}}$ and $\nabla_{t}\mathcal{L}_{\|\nabla\mathcal{L}_{b}^{\perp}}$ are used to represent $\nabla\mathcal{L}(\theta_{t})_{\|(\nabla\mathcal{L}_{r}(\theta_{t}))^{\perp}}$ and $\nabla_{t}\mathcal{L}(\theta_{t})_{\|(\nabla\mathcal{L}_{b}(\theta_{t}))^{\perp}}$, the main difference is the way of picking a common descent direction that belongs to $\mathbf{G}_{t}$. The pseudo code of algorithms DCGD-Center, DCGD-Average, and DCGD-Proj are shown as follows:
1 Require: learning rate $\lambda$, max epoch $T$, initial point $\theta_{0}$, gradient threshold $\varepsilon$, conflict threshold $\alpha$

### v3

@@ Abstract
Physics-informed machine learning (PIML) has been widely adopted for traffic flow modeling in recent studies, due to its potential in combining the benefits of both physics-based and data-driven approaches. In conventional PIML, physics information from classical traffic flow models is typically incorporated by constructing a hybrid loss function that combines data-driven loss and physics loss through linear scalarization. The goal is to find a trade-off between these two objectives to improve the accuracy of model predictions. However, from a mathematical perspective, linear scalarization is limited to identifying only the convex region of the Pareto front, as it treats data-driven and physics losses as separate objectives. Given that most PIML loss functions are non-convex, linear scalarization restricts the achievable trade-off solutions. Moreover, tuning the weighting coefficients for the two loss components can be both time-consuming and computationally challenging. To address these limitations, this paper introduces a paradigm shift in PIML by reformulating the training process as a multi-objective optimization problem, treating data-driven loss and physics loss independently. We apply several multi-gradient descent algorithms (MGDAs), including traditional multi-gradient descent (TMGD) and dual cone gradient descent (DCGD), to explore the Pareto front in this multi-objective setting. These methods are evaluated on both macroscopic and microscopic traffic flow models. In the macroscopic case, MGDAs achieved comparable performance to traditional linear scalarization methods. Notably, in the microscopic case, MGDAs significantly outperformed their scalarization-based counterparts, demonstrating the advantages of a multi-objective optimization approach in complex PIML scenarios.
@@ 1 Introduction
PIML effectively combines the advantages of data-driven and physics-based methods, making it useful in transportation research. Data-driven approaches have gained popularity due to their low computational cost and flexibility in managing complex scenarios without requiring strong theoretical assumptions. Many of machine learning techniques, especially deep learning architectures Ma et al. [2015], Duan et al. [2016], Polson and Sokolov [2017], Zhou et al. [2017], Wu et al. [2018], Cui et al. [2018, 2019], Zhang et al. [2025], have various application in transportation research. However, these models heavily depend on the quality and representativeness of the training data. As shown in Yuan et al. [2021b], when manually adding errors to the training dataset, the performance of all pure data-driven models has significantly deteriorated. While a data-driven approach can yield positive results with quality data, it often lacks interpretability, which may diminish its practicality in decision-making processes Lei et al. [2024].
@@
In nearly all PIML studies, especially in PIML application in transportation, as shown in Table 1, researchers commonly use a method called linear scalarization to define the loss function [^1] In many Physics-regularized Gaussian process models (Yuan et al. [2020],Yuan et al. [2021b],Yuan et al. [2021a]), the training target is to minimize a evidence lower bound, which can be written as: $\log\mathcal{N}([\cdot])+\gamma\mathbb{E}(\cdot)$ (1) where $\gamma$ is a regularization coefficient. The overall ELBO can also be regarded as linear scalarization., as demonstrated in Eq 1. where $\alpha$ and $\beta$ are hyperparameters that control the effect of the data-driven module and physics module, $\mathcal{L}_{data}$ and $\mathcal{L}_{physics}$ represent data-driven loss term and physics loss term, respectively. This approach facilitates the easy integration of physics information into PIML models. When $\alpha\rightarrow 0$, the PIML model degenerates to a pure data-driven model; when $\beta\rightarrow 0$, the PIML model degenerates to a pure physics-driven model.
@@
In this paper, we aim to circumvent the hyperparameter search process by transforming the training process of the PIML model into a multi-objective optimization problem. Mathematically, we transfer the overall training process of PIML models from Eq 3 to 4. To clarify, in the rest of the paper, multi-objective optimization refers to the core training paradigm we adopt, which treats the data loss and physics loss as distinct objectives and optimizes them simultaneously.
@@ 2 Literature review
One clear strength of the PIML framework is its flexibility: the learning module can adopt almost any neural-network architecture—or even non-NN techniques—to suit a particular task. For instance, Xue et al. [2024] proposes a Graph Neural Network (GNN) with the Network Macroscopic Fundamental Diagram (NMFD) to impute missing traffic states. The GNN captures complex spatial dependencies among links, while the NMFD term keeps predictions physically plausible. Likewise, Pereira et al. [2022] shows that a long short-term memory (LSTM) backbone plugs in just as naturally, enabling short-range traffic forecasting when strong temporal correlations exist. In both cases, the physics term reduces data needs and stabilizes training.
Besides using Neural Networks, Gaussian Process has also been a widely-adopted tool for traffic flow modeling. Xu et al. [2023] proposed a method to integrate attentive graph neural process (AGNP) and Gaussian process for network-level short-term traffic speed prediction and imputation. Zhu et al. [2023] develops a Bayesian clustering ensemble Gaussian process (BCEGP) model for network-wide traffic flow clustering and prediction. Integrated with physics knowledge and inspired by Wang et al. [2020], Yuan et al. [2021b] introduce the Physics-Regulated Gaussian Process (PRGP). Here, the governing equations appear through a shadow Gaussian process rather than an extra PINN loss term. Training maximizes a compound evidence lower bound (ELBO) that balances the usual GP likelihood with a penalty that measures how well samples obey the flow equation. A follow-up study by Yuan et al. [2021a] extends this idea to a richer family of macroscopic models, improving fidelity and robustness without sacrificing the analytic uncertainty estimates that make Gaussian processes attractive.
@@ 3.1 Multi-gradient descent algorithms
    There exist $\alpha^{1},...,\alpha^{T}$ such that $\sum_{t=1}^{T}\alpha^{t}=1$ and $\sum_{t=1}^{T}\alpha^{t}\nabla_{\boldsymbol{\theta}^{sh}}\mathcal{L}^{t}(\boldsymbol{\theta}^{sh},\boldsymbol{\theta}^{t})=0$
@@
    For all objectives $t$, $\nabla_{\boldsymbol{\theta}^{t}}\mathcal{L}^{t}(\boldsymbol{\theta}^{sh},\boldsymbol{\theta}^{t})=0$
where $\boldsymbol{\theta}^{sh}$ is the shared parameters among all objectives, $\boldsymbol{\theta}^{t}$ is the exclusive parameters of $t$th objective. [^2] In the context of PIML, $\boldsymbol{\theta}^{sh}\subset\boldsymbol{\theta}^{data}\cap\boldsymbol{\theta}^{physics}$
In addition, readers can find out that some of the studies about multi-gradient descent (Hwang and Lim [2024]) do not divide the total parameter set $\boldsymbol{\theta}$ into $\boldsymbol{\theta}^{sh}$, $\boldsymbol{\theta}^{data}$ and $\boldsymbol{\theta}^{physics}$ like in Sener and Koltun [2018]. In fact, let us consider the data-driven loss $\mathcal{L}_{data}$, if we calculate its derivatives of the entire parameter set $\boldsymbol{\theta}$, we will have:
@@ 3.1.1 Traditional multi-gradient descent algorithm
The convex hull of a set of points $S$ in $n$ dimensions is defined as the intersection of all convex sets containing $S$. For $N$ points $\mathbf{p}_{1}$, $\mathbf{p}_{2}$$\mathbf{p}_{N}$, the convex hull $\textbf{conv}(\mathbf{p}_{1},\mathbf{p}_{2},..,\mathbf{p}_{N})$ is then given by:
@@
Let $C\subset\mathbb{R}^{n}$ be a nonempty closed convex set. The minimum norm point of $C$ is the unique solution
@@
Let $P$ $\subset$ $\mathbb{R}^{d}$ be the convex hull of finitely many points, $P$ = $\mathbf{conv}(\mathbf{p}_{1}$, $\mathbf{p}_{2}$$\mathbf{p}_{n}$). A point $\mathbf{x}\in P$ is the minimum norm point in $P$ if and only if:
@@
Let $P$ $\subset$ $\mathbb{R}^{d}$ be the convex hull of the gradients of the different objective functions, i.e., $P$ = conv($\mathbf{p}_{1}$, $\mathbf{p}_{2}$$\mathbf{p}_{n}$). If $\mathbf{x}$ is the minimum norm point, and $\mathbf{x}\in P$, then $\mathbf{x}$ is a direction that can improve all objectives.
@@
[…]
Recall that $\mathbf{p}_{j}$ is the gradient with respect to $j$th objective function. From (10), the directional derivative of each objective in the direction $\mathbf{x}$ is non-positive. Therefore, $\mathbf{x}$ constitutes a direction that does not increase (and thus improve) all objectives. In particular, when $\mathbf{x}$ $\neq$ $\mathbf{0}$, $\mathbf{x}$ strictly decreases (and thus improves) all objectives, completing the proof. ∎
@@ 10     $\hat{t}=\operatorname*{arg\,min}_{r}\sum_{t}\alpha^{t}\mathbf{M}_{rt}$
[…]
@@
Either solution to this optimization problem is 0 and the resulting point satisfies the KKT conditions, or the solution gives a descent direction $\sum_{j=1}^{N}\alpha_{j}^{*}\nabla_{\boldsymbol{\theta}^{sh}}{\mathcal{F}}^{\top}(\boldsymbol{\boldsymbol{\theta}^{sh},\boldsymbol{\theta}^{j}})$ that improves all objectives.
@@ 3.1.2 Dual cone gradient descent
In this section, $\nabla\mathcal{L}_{p}(\boldsymbol{\theta}_{t})$, $\nabla\mathcal{L}_{d}(\boldsymbol{\theta}_{t})$, $\nabla\mathcal{L}(\boldsymbol{\theta}_{t})$ represent physics loss, data loss, and total loss, respectively. $\boldsymbol{\theta}_{t}$ is the overall parameter set for all physics and data loss in $t$th iteration. $\phi_{t}$ is the angle between $\nabla\mathcal{L}_{d}(\boldsymbol{\theta}_{t})$ and $\nabla\mathcal{L}_{p}(\boldsymbol{\theta}_{t})$. $\|$ is the parallel component symbol, the parallel component of the vector $\mathbf{v}$ along $\mathbf{a}$, defined by:
@@
The concept of dual cone shown in Definition 6 plays an important role in the DCGD algorithm. However, in practice, we are more interesting of its subspace $\mathbf{G}_{t}$, as shown in Figure 1 and mathematically defined in Eq 15. Any direction such that $\mathbf{d}\in\mathbf{G}_{t}$ is a common descent direction since $\langle\mathbf{d},\nabla\mathcal{L}_{r}(\theta_{t})\rangle\geq 0$ and $\langle\mathbf{d},\nabla\mathcal{L}_{b}(\theta_{t})\rangle\geq 0$. Also, from Eq 15, $\mathbf{G}_{t}$ is essentially a conic combination of $\nabla\mathcal{L}(\theta_{t})_{\|(\nabla\mathcal{L}_{p}(\theta_{t}))^{\perp}}$ and $\nabla_{t}\mathcal{L}(\theta_{t})_{\|(\nabla\mathcal{L}_{d}(\theta_{t}))^{\perp}}$, which is very easy to compute compare to compute the $\mathbf{K}^{*}_{t}$.
@@
$\mathbf{G}_{t}:=\left\{c_{1}\nabla\mathcal{L}(\theta_{t})_{\|(\nabla\mathcal{L}_{p}(\theta_{t}))^{\perp}}+c_{2}\nabla_{t}\mathcal{L}(\theta_{t})_{\|(\nabla\mathcal{L}_{d}(\theta_{t}))^{\perp}}\big|c_{1},c_{2}\geq 0,c_{1},c_{2}\in\mathbb{R}\right\}$ (15)
@@
Algorithm 2 shows the core idea of DCGD, that is to find a common descent direction from $\mathbf{G}^{*}_{t}$, and since any direction that subset to $\mathbf{G}_{t}$ is a valid common descent direction. Thus, as shown in Figure 2, there are several variants of DVGD algorithms [^3] For notation simplicity, $\nabla\mathcal{L}_{\|\nabla\mathcal{L}_{p}^{\perp}}$ and $\nabla_{t}\mathcal{L}_{\|\nabla\mathcal{L}_{d}^{\perp}}$ are used to represent $\nabla\mathcal{L}(\theta_{t})_{\|(\nabla\mathcal{L}_{p}(\theta_{t}))^{\perp}}$ and $\nabla_{t}\mathcal{L}(\theta_{t})_{\|(\nabla\mathcal{L}_{d}(\theta_{t}))^{\perp}}$, the main difference is the way of picking a common descent direction that belongs to $\mathbf{G}_{t}$. The pseudo code of algorithms DCGD-Center, DCGD-Average, and DCGD-Proj are shown as follows:
1Require: learning rate $\lambda$, max epoch $T$, initial point $\theta_{0}$, gradient threshold $\varepsilon$, conflict threshold $\alpha$
2 for $t=1$ to $T$ do
3   if $\pi-\alpha<\phi_{t}\text{ or }\|\nabla\mathcal{L}(\boldsymbol{\theta}_{t})\|<\varepsilon$ then
4     break
5   end if
6    $g_{t}^{c}=\frac{\nabla\mathcal{L}_{d}(\boldsymbol{\theta}_{t})}{\|\nabla\mathcal{L}_{d}(\boldsymbol{\theta}_{t})\|}+\frac{\nabla\mathcal{L}_{p}(\boldsymbol{\theta}_{t})}{\|\nabla\mathcal{L}_{p}(\boldsymbol{\theta}_{t})\|}$
7    $g_{t}^{dual}=\frac{\langle g_{t}^{c},\nabla\mathcal{L}(\boldsymbol{\theta}_{t})\rangle}{\|g_{t}^{c}\|^{2}}g_{t}^{c}$
8    $\boldsymbol{\theta}_{t}=\boldsymbol{\theta}_{t-1}-\lambda g_{t}^{dual}$
9 end for
Algorithm 2 DCGD (Center) [Hwang and Lim [2024]]
1Require: learning rate $\lambda$, max epoch $T$, initial point $\theta_{0}$, gradient threshold $\varepsilon$, conflict threshold $\alpha$